A writer received $37,000 as royalty for her book. She invested part of the money in bonds paying 6% interest annually. The rest she invested in a life insurance policy paying 11% interest annually. If the total interest from the investments after 1 year is $3520 how much did she invest in bonds?
Answer by mananth(16946)

Investment I 6.00% per annum ---x
Investment II 11.00% per annum ---y

x+y=37000 ------------------------1
6.00%x+ 11.00% y=$3,520.00
Multiply by 100
6x+11y= $352,000.00 --------2

Multiply (1) by -6
we get

-6x-6y= -222000.00

Add this to (2)

0 x+ 5y= =$130,000.00

divide by 5

y =$26,000.00 investment at 11.00%
Balance $11,000.00 investment at 6.00%
CHECK
$11,000.00 --------- 6.00% ------- $660.00
$26,000.00 ------- 11.00% ------- $2,860.00
Total -------------- $3,520.00
m.ananth@hotmail.ca